FIND AREA OF THE GRAPH UPPER AND LOWER SUMS?!?!?

Use upper and lower sums to approximate the area of the region using the given number of subintervals (of equal width). (Round your answers to three decimalplaces.)
y = 4e−x

upper sum__________

lower sum___________


Thanks for the help!!! :)
Answer #1
4.3
2.1
answered by: noob
Answer #2

y = sqrt (1-x^2)

Each sub interval needs to be 0.2 units

Upper sums
y(0) * 0.2 + y(0.2)*0.2 + y(0.4) *0.2 + y(0.6) * 0.2 + y(0.8) * 0.2
= (1/5)(y(0) + y(0.2) + y(0.4) + y(0.6) + y(0.8))

Lower sums
= (1/5)(y(0.2) + y(0.4) + y(0.6) + y(0.8) + y(1))

y(0) = 1
y(0.2) = sqrt (1-1/25) = sqrt (24/25)
y(0.4) = sqrt (1-4/25) = sqrt (21/25)
y(0.6) = sqrt (1-9/25) = sqrt (16/25) = 4/5
y(0.8) = sqrt (1-16/25) = sqrt (9/25) = 3/5
y(1) = 0
Upper sums =
0.8592622 units^2 (7dp)
Lower sums =
0.6592622 units^2 (7dp)

If you want to average these
A = 0.7592622 units^2 (7dp)


Since its 1/4 of a circle with radius = 1
Area = pi*162 / 4
=0.7853981634 units^2 (10dp)

answered by: mkcolema